if an equilateral triangle has a perimeter of (2x+9)miles and one side of the triangle measures (x+2) miles, how long (in miles)is the side of the triangle ?
a 3 b 5 c 9 d 15

Answer by mananth(16949)

perimeter = 2x+9
one side = x+2
3*(x+2)=2x+9
3x+6=2x+9
3x-2x=9-6
x=3
one side = x+2 = 3+2=5
side =5